# Autoregressive Language Models are Secretly Energy-Based Models: Insights into the Lookahead Capabilities of Next-Token Prediction

### v1

@@ Abstract
Autoregressive models (ARMs) currently constitute the dominant paradigm for large language models (LLMs). Energy-based models (EBMs) represent another class of models, which have historically been less prevalent in LLM development, yet naturally characterize the optimal policy in post-training alignment. In this paper, we provide a unified view of these two model classes. Taking the chain rule of probability as a starting point, we establish an explicit bijection between ARMs and EBMs in function space, which we show to correspond to a special case of the soft Bellman equation in maximum entropy reinforcement learning. Building upon this bijection, we derive the equivalence between supervised learning of ARMs and EBMs. Furthermore, we analyze the distillation of EBMs into ARMs by providing theoretical error bounds. Our results provide insights into the ability of ARMs to plan ahead, despite being based on the next-token prediction paradigm.
@@ 2.2 Autoregressive models (ARMs)
where ${\mathcal{S}}({\bm{x}})\subseteq{\mathcal{S}}$ is the set of partial sequences prolonging the prompt $\bm{x}$}. We assume that we have at our disposal a function $q\colon{\mathcal{S}}\times{\mathcal{A}}$, for scoring the “affinity” between a context (state) ${\bm{s}}_{t}\in{\mathcal{S}}$ and a next token $y_{t}\in{\mathcal{A}}$ (potentially EOS). We can think of $q$ as a local scoring function, while $R$ used by EBMs is a global scoring function. ARMs use the chain rule of probability to define the probability of a sequence $\bm{y}$} given $\bm{x}$} as
@@ 3.1 Bijection between distributions
3.2 Bijection between EBMs and ARMs
@@
where $r\colon{\mathcal{S}}\times{\mathcal{A}}\to{\mathbb{R}}$ is the immediate (next-token) reward. Indeed, if $R$ does not naturally decompose as above, from any $R\colon{\mathcal{X}}\times{\mathcal{Y}}\to{\mathbb{R}}$ , we can always define
$r({\bm{s}}_{t},y_{t})\coloneqq\begin{cases}0&\mbox{if}~y_{t}\neq\texttt{EOS}\\ R({\bm{x}},{\bm{y}})&\mbox{if}~y_{t}=\texttt{EOS}\\ \end{cases}.$ (21)
@@
where ${\bm{s}}_{t}\in{\mathcal{S}}$ and $y_{t}\in{\mathcal{A}}$. For all $\tau$ and all ${\bm{y}}=(y_{1},\dots,y_{\tau})$, defining the contexts ${\bm{s}}_{1}\coloneqq{\bm{x}}$, …, ${\bm{s}}_{\tau}\coloneqq{\bm{x}}\oplus{\bm{y}}_{<\tau}$ we obtain by unrolling (22) backwards,
$q({\bm{s}}_{\tau},y_{\tau})$ $=r({\bm{s}}_{\tau},y_{\tau})$ (23)
$\vdots$ (24)
$q({\bm{s}}_{t},y_{t})$ $=r({\bm{s}}_{t},y_{t})+V_{q}({\bm{s}}_{t}\oplus y_{t})$ (25)
$\vdots$ (26)
$q({\bm{s}}_{1},y_{1})$ $=r({\bm{s}}_{1},y_{1})+V_{q}({\bm{s}}_{1}\oplus y_{1}).$ (27)
Because $V_{q}({\bm{s}}_{t}\oplus y_{t})=\underset{y_{t+1}\in{\mathcal{A}}}{{\mathrm{LSE}}}~q({\bm{s}}_{t}\oplus y_{t},y_{t+1})$, we can see that $q({\bm{s}}_{t},\cdot)$ depends on $q({\bm{s}}_{t+1},\cdot)$ and therefore computing $q$ from $r$ must be done sequentially for all possible states.
@@
Applying (28), we obtain
$r({\bm{s}}_{1},y_{1})$ $=q({\bm{s}}_{1},y_{1})-V_{q}({\bm{s}}_{1}\oplus y_{1})$ (29)
$\vdots$ (30)
$r({\bm{s}}_{t},y_{t})$ $=q({\bm{s}}_{t},y_{t})-V_{q}({\bm{s}}_{t}\oplus y_{t})$ (31)
$\vdots$ (32)
$r({\bm{s}}_{\tau},y_{\tau})$ $=q({\bm{s}}_{\tau},y_{\tau}).$ (33)
Computing $r$ from $q$ can be done in parallel.
@@
Having defined the mapping $\mathcal{M}$} and its inverse ${\mathcal{M}}^{-1}$, we can now formally state our bijection.
@@
[Bijection between EBMs and ARMs] If $q={\mathcal{M}}(r)$ or equivalently $r={\mathcal{M}}^{-1}(q)$, then, for all ${\bm{x}}\in{\mathcal{X}}$ and ${\bm{y}}\in{\mathcal{Y}}$, we have
@@
The cost of converting an EBM into an ARM is $O(V^{T})$, exponential in the sequence length. This is due to the fact that the log-sum-exp on time $t$ depends on $q({\bm{s}}_{t+1},\cdot)$, which creates a recursive dependency. The cost of converting an ARM to an EBM, however, is $O(VT)$, linear in the sequence length. Again, the log-sum-exp on time $t$ depends on $q({\bm{s}}_{t+1},\cdot)$. The crucial difference is that, this time, $q$ is given, not recursively defined. This makes it possible to compute $r({\bm{s}}_{t},y_{t})$ for all $t\in[T]$, in parallel.
@@ 3.4 Entropy-regularized DP on a DAG perspective
$=\underset{y_{1}\in{\mathcal{V}}}{{\mathrm{LSE}}}~\biggl[r({\bm{x}},y_{1})+\underset{y_{2}\in{\mathcal{V}}}{{\mathrm{LSE}}}~\Bigl[r({\bm{x}}\oplus y_{1},y_{2})+\dots$ (43)
$+\underset{y_{T}\in{\mathcal{V}}}{{\mathrm{LSE}}}~r({\bm{x}}\oplus{\bm{y}}_{<T},y_{T})\Bigr]\biggr].$ (44)
@@ 4.1 Learning EBMs and ARMs from input-output pairs
This demonstrates that teacher forcing is optimal when optimizing in function space. We emphasize that the mapping $\mathcal{M}$} is completely implicit in this case (we do not need to explicitly compute it).
@@ 4.2 Distilling EBMs into ARMs
In practice, $q$ and $r$ are implemented using parameterized functions. In this setting, there may no longer be an exact mapping between EBMs and ARMs, but we can bound their KL divergence. {proposition}[KL bound] For all $r\colon{\mathcal{S}}\times{\mathcal{A}}\to{\mathbb{R}}$, $q\colon{\mathcal{S}}\times{\mathcal{A}}\to{\mathbb{R}}$ and ${\bm{x}}\in{\mathcal{X}}$, we have
@@ 6 Related work
Maxent RL can be viewed from a probabilistic inference perspective, using undirected graphical models (Toussaint et al., 2009; Ziebart, 2010), defining EBMs (a.k.a. conditional random fields), or directed graphical models, mirroring hidden Markov models (Attias, 2003; Levine, 2018). Notably, this connection has been leveraged to solve inference problems using MCTS and learned soft value functions (Buesing et al., 2020). Conversely, probabilistic inference tools can be leveraged for maxent RL (Korbak et al., 2022), such as twisted sequential Monte Carlo (Zhao et al., 2024). The equivalence between optimal estimation (Kalman filtering) and optimal control (for linear-quadratic systems) is known as Kalman duality (Todorov, 2008). We revisit the probabilistic inference perspective in our setting and notation in Appendix A.9.
@@
The standard approach for learning an EBM from $({\bm{x}},{\bm{y}})$ pairs by stochastic gradient descent requires sampling from the EBM (Song & Kingma, 2021), which usually involves MCMC algorithms such as Langevin when $\mathcal{Y}$} is continuous or Gibbs sampling when $\mathcal{Y}$} is discrete. Since that approach learns the negative energy $R$ (the reward in inverse RL; see Section 4.1), and not a generative model $\pi$, MCMC is needed for generating samples from the EBM parameterized by the learned $R$. A min-min formulation was proposed to jointly learn $R$ and the log-partition $A$ (Sander et al., 2025). While this bypasses the need for MCMC at train time, it still requires it for sampling from the EBM. We emphasize however that if we do not need to sample but just to predict the mode of the EBM (the most likely output $\bm{y}$} given $\bm{x}$}), then we only need to maximize $R$, which does not involve the log-partition and is tractable for some sets $\mathcal{Y}$}. On the other hand, min-max formulations can be used to jointly learn a discriminator (reward) $R$ and a generator (policy) $\pi$ from $({\bm{x}},{\bm{y}})$ pairs (Ho & Ermon, 2016). Such approach does not require MCMC if the generator easily generates samples by design, as is the case with ARMs.
@@
GFlowNets (Bengio et al., 2021, 2023) can be used to learn discrete samplers for amortized probabilistic inference. Flow consistency equations play the same role as soft Bellman equations, enabling the design of consistency loss functions (Malkin et al., 2022). In fact, recent work has established that GFlowNets can be cast as MaxEnt RL, even in the general DAG setting (Tiapkin et al., 2024).
@@
Recent work argued that the next-token prediction objective (teacher forcing) is structurally “myopic” and fails to backpropagate future value information effectively, leading to a failure in planning (Bachmann & Nagarajan, 2024). In contrast, Proposition 4.1 shows that the global minimum of the teacher forcing objective (in function space) is an EBM, which inherently performs lookahead via the partition function $V_{q}$. Our work therefore suggests that the limitation of teacher forcing is not in the objective itself but in the difficulty of reaching its minimum, due to optimization or approximation error. This aligns with prior work showing that planning failures stem from overfitting to spurious correlations under weak supervision, rather than inherent structural incapacity (Frydenlund, 2024).
@@ B.1 Lemmas
Because we can always define $r$ from $R$ using (21), we have
@@ C.3 Metrics illustrated
To illustrate Proposition 4.1, we plot the losses along the training of these Transformers. Namely, in the left panel of Figures 12 and 3, we plot
${\mathcal{L}}^{\textsc{ebm}}(R_{w_{k}})-{\mathcal{L}}^{\star},\quad\text{and}\quad{\mathcal{L}}^{\textsc{arm}}(q_{w_{k}})-{\mathcal{L}}^{\star},$
@@
where $c({\bm{z}})={\bm{z}}-\langle{\bm{z}},{\mathbf{1}}\rangle{\mathbf{1}}$. In Figures 12 and 3, we call this measurement of the logits distance “Before mapping”.
@@
where $c_{-1}$ denotes centering along the last axis. In Figures 12 and 3, we call this measurement the logits distance “After mapping”.

### v2

@@ Abstract
Autoregressive models (ARMs) currently constitute the dominant paradigm for large language models (LLMs). Energy-based models (EBMs) represent another class of models, which have historically been less prevalent in LLM development, yet naturally characterize the optimal policy in post-training alignment. In this paper, we present a unified view of these two model classes. Taking the chain rule of probability as a starting point, we establish an explicit bijection between ARMs and EBMs in function space, which we show to correspond to a special case of the soft Bellman equation in maximum entropy reinforcement learning. Building upon this bijection, we derive the equivalence between supervised learning of ARMs and EBMs. Furthermore, we analyze the distillation of EBMs into ARMs by providing theoretical error bounds. Our results provide insights into the ability of ARMs to plan ahead, despite being based on the next-token prediction paradigm.
@@ 2.2 Autoregressive models (ARMs)
where ${\mathcal{S}}({\bm{x}})\subseteq{\mathcal{S}}$ is the set of partial sequences prolonging the prompt $\bm{x}$}. We assume that we have at our disposal a function $q\colon{\mathcal{S}}\times{\mathcal{A}}$, for scoring the “affinity” between a context (state) ${\bm{s}}_{t}\in{\mathcal{S}}$ and a next token $y_{t}\in{\mathcal{A}}$ (potentially EOS). We can think of $q$ as a local scoring function, while $R$ used by EBMs is a global scoring function. ARMs (Bengio et al., 2003) use the chain rule of probability to define the probability of a sequence $\bm{y}$} given $\bm{x}$} as
@@
Figure 1: Summary of mappings discussed in this paper.
@@ 3.1 Bijection between distributions
3.2 Bijection in function space
@@
where $r\colon{\mathcal{S}}\times{\mathcal{A}}\to{\mathbb{R}}$ is the immediate (next-token) reward. Indeed, if $R$ does not naturally decompose as above, we can always assign all the non-zero reward in the last step.
@@
where ${\bm{s}}_{t}\in{\mathcal{S}}$ and $y_{t}\in{\mathcal{A}}$. Recall that
$V_{q}({\bm{s}}_{t}\oplus y_{t})=\underset{y_{t+1}\in{\mathcal{A}}}{{\mathrm{LSE}}}~q({\bm{s}}_{t}\oplus y_{t},y_{t+1}).$ (22)
We can see that $q({\bm{s}}_{t},\cdot)$ depends on $q({\bm{s}}_{t+1},\cdot)$, creating a recursive dependency. Computing $q$ from $r$ must thus be done sequentially from $t=\tau$ to $t=1$ for all possible states.
@@
As there is no longer a recursive dependency, computing $r$ from $q$ can be done in parallel.
@@
Having defined the mapping $\mathcal{M}$} and its inverse ${\mathcal{M}}^{-1}$, we can now formally state our result.
@@
[Bijection in function space] The mapping $q={\mathcal{M}}(r)$ is bijective and for all ${\bm{x}}\in{\mathcal{X}}$ and ${\bm{y}}\in{\mathcal{Y}}$, we have
@@
The cost of converting an ARM to an EBM is $O(VT)$, linear in the sequence length. On the other hand, the cost of converting an EBM into an ARM is $O(V^{T})$, exponential in the sequence length. Explicitly converting an EBM into an ARM is therefore intractable. However, as we shall discuss in Section 4.1, when learning from $({\bm{x}},{\bm{y}})$ pairs, such an explicit conversion is not needed, as the optimal ARM is implicitly an EBM!
@@ 3.4 Entropy-regularized DP on a DAG perspective
$=\underset{y_{1}\in{\mathcal{V}}}{{\mathrm{LSE}}}~\biggl[r({\bm{x}},y_{1})+\dots+\underset{y_{T}\in{\mathcal{V}}}{{\mathrm{LSE}}}~\Bigl[r({\bm{x}}\oplus{\bm{y}}_{<T},y_{T})\Bigr]\biggr].$ (33)
@@ 4.1 Learning EBMs and ARMs from input-output pairs
This demonstrates that teacher forcing is optimal when optimizing in function space. We stress that the mapping $\mathcal{M}$} is completely implicit in this case (we do not need to explicitly compute it). We summarize the mappings in Figure 1.
@@ 4.2 Distilling EBMs into ARMs
In practice, $q$ and $r$ are implemented using parameterized functions. In this setting, there may no longer be an exact conversion between EBMs and ARMs, but we can bound their KL divergence. {proposition}[KL bound] For all $r\colon{\mathcal{S}}\times{\mathcal{A}}\to{\mathbb{R}}$, $q\colon{\mathcal{S}}\times{\mathcal{A}}\to{\mathbb{R}}$ and ${\bm{x}}\in{\mathcal{X}}$, we have
@@ 6 Related work
Maxent RL can be viewed from a probabilistic inference perspective, using undirected graphical models (Toussaint et al., 2009; Ziebart, 2010), defining EBMs (a.k.a. conditional random fields), or directed graphical models, mirroring hidden Markov models (Attias, 2003; Levine, 2018). Notably, this connection has been leveraged to solve inference problems using MCTS and learned soft value functions (Buesing et al., 2020). Conversely, probabilistic inference tools can be leveraged for maxent RL (Korbak et al., 2022), such as twisted sequential Monte Carlo (Zhao et al., 2024). We revisit the probabilistic inference perspective in our setting in Appendix A.9.
@@
The standard approach for learning an EBM from $({\bm{x}},{\bm{y}})$ pairs by stochastic gradient descent requires sampling from the EBM (Song & Kingma, 2021), which usually involves MCMC algorithms such as Langevin when $\mathcal{Y}$} is continuous or Gibbs sampling when $\mathcal{Y}$} is discrete. Since that approach learns the function $R$ (the reward in inverse RL; see Section 4.1), and not a generative model $\pi$, MCMC is needed for generating samples from the EBM parameterized by the learned $R$. A min-min formulation was proposed to jointly learn $R$ and the log-partition $A$ (Sander et al., 2025). While this bypasses the need for MCMC at train time, it still requires it for sampling from the EBM. We emphasize however that if we do not need to sample but just to predict the mode of the EBM (the most likely output $\bm{y}$} given $\bm{x}$}), then we only need to maximize $R$, which does not involve the log-partition and is tractable for some sets $\mathcal{Y}$}. On the other hand, min-max formulations can be used to jointly learn a discriminator (reward) $R$ and a generator (policy) $\pi$ from $({\bm{x}},{\bm{y}})$ pairs (Ho & Ermon, 2016). Such approach does not require MCMC if the generator easily generates samples by design, as is the case with ARMs.
@@
GFlowNets (Bengio et al., 2021, 2023) can be used to perform amortized probabilistic inference (Hu et al., 2023). Flow consistency equations play the same role as soft Bellman equations, enabling the design of consistency loss functions (Malkin et al., 2022). In fact, recent work has established that GFlowNets can be cast as MaxEnt RL, even in the general DAG setting (Tiapkin et al., 2024).
@@
Recent work argued that the next-token prediction objective (teacher forcing) is structurally “myopic” and fails to backpropagate future value information effectively, leading to a failure in planning (Bachmann & Nagarajan, 2024). In contrast, Proposition 4.1 shows that the global minimum of the teacher forcing objective (in function space) is an EBM, which inherently performs lookahead via the partition function $V_{q}$. Our work therefore suggests that the limitation of teacher forcing is not in the objective itself but in the difficulty of reaching its minimum, due to optimization or approximation error. This aligns with prior work showing that planning failures stem from overfitting to spurious correlations under weak supervision, rather than inherent structural incapacity (Frydenlund, 2024). In complementary work, behavior cloning, the counterpart of teacher forcing in RL, was shown to achieve horizon-independent sample complexity (Foster et al., 2024).
@@ 7 Conclusion
Impact statement
This paper studies the connection between energy-based models and autoregressive models to justify the lookahead capabilities of next-token prediction. We do not foresee any specific ethical or societal implications arising directly from this work.
@@ B.1 Lemmas
The existence of the explicit inverse ${\mathcal{M}}^{-1}$ ensures that $\mathcal{M}$} is a bijective mapping.
@@
Because we can always define $r$ from $R$ using
$r({\bm{s}}_{t},y_{t})\coloneqq\begin{cases}0&\mbox{if}~y_{t}\neq\texttt{EOS}\\ R({\bm{x}},{\bm{y}})&\mbox{if}~y_{t}=\texttt{EOS}\\ \end{cases},$ (182)
we have
@@ C.3 Metrics illustrated
To illustrate Proposition 4.1, we plot the losses along the training of these Transformers. Namely, in the left panel of Figures 23 and 4, we plot
${\mathcal{L}}^{\textsc{ebm}}(R(w_{k}))-{\mathcal{L}}^{\star},\quad\text{and}\quad{\mathcal{L}}^{\textsc{arm}}(q(w_{k}))-{\mathcal{L}}^{\star},$
@@
where $c({\bm{z}})={\bm{z}}-\langle{\bm{z}},{\mathbf{1}}\rangle{\mathbf{1}}$. In Figures 23 and 4, we call this measurement of the logits distance “Before mapping”.
@@
where $c_{-1}$ denotes centering along the last axis. In Figures 23 and 4, we call this measurement the logits distance “After mapping”.